Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Contidional Probability?

Expert replies
by sumithshah » Sat Sep 27, 2008 11:13 pm
If Ben were to lose the championship, Mike would be the winner with a probability of 1/4 , and Rob 1/3 . If the probability of Ben being the winner is , what is the probability that either Mike or Rob will win the championship?

1/12 ; 1/7 ; 1/2 ; 7/12 ; 6/7

OA C


I have no clue how do this. Any pointers would be appreciated
Last edited by sumithshah on Sun Sep 28, 2008 12:24 am, edited 1 time in total.
Join the discussion
Source: — Problem Solving |

by manulath » Sat Sep 27, 2008 11:58 pm
please check the question, it is missing data - prababilities!
regards
Join the discussion

by sumithshah » Sun Sep 28, 2008 12:24 am
fixed
Join the discussion

by 4meonly » Sun Sep 28, 2008 2:37 am
not fixed yet

winner is ???????, what is the
Join the discussion

Re: Contidional Probability?

by manulath » Sun Sep 28, 2008 3:17 am
[quote="sumithshah"]If the probability of Ben being the winner is ,


Please correct the question. Data missing.
Join the discussion

Re: Contidional Probability?

by sudhir3127 » Sun Sep 28, 2008 3:27 am
sumithshah wrote:If Ben were to lose the championship, Mike would be the winner with a probability of 1/4 , and Rob 1/3 . If the probability of Ben being the winner is , what is the probability that either Mike or Rob will win the championship?

1/12 ; 1/7 ; 1/2 ; 7/12 ; 6/7

OA C


I have no clue how do this. Any pointers would be appreciated


This is the right Question

"If Ben were to lose the championship, Mike would be the winner with a probability of 1/4, and Rob - 1/3. If the probability of Ben being the winner is 1/7, what is the probability that either Mike or Rob will win the championship?" Sumit would appreciate if u double check before posting a Question .. it would unnecessarily delay the posters time ....


here goes the Answer...

Prob of Ben wins=1/7

Prob of ben loses =6/7.

P(Mike wins or Rob wins) = P(Mike wins)+P(Rob wins) - P(Both wins)
=1/4+1/3-0
=7/12.

But we are only 6/7 sure about Ben loosing.

hence the probability is = 6/7*7/12=1/2

hope that helps....
Join the discussion

by uptowngirl92 » Fri Sep 18, 2009 5:30 pm
P(Mike winning)=P(Ben looses). P(Mike wins) . P(Rob looses)
P(Rob winning)=P(Ben looses). P(Mike looses) . P(Rob wins)

Adding the two will give us the answer right?

Now given,P(mike wins)=1/4.Therefore,P(mike looses)=3/4
Similarly,P(rob wins)=1/3.Therefore,P(rob looses)=2/3
And,P(ben wins)=1/7.Therefore,P(ben looses)=6/7

Putting the values,
6/7.1/4.2/3 + 6/7.3/4.1/3 = 1/7+3/14=5/14 :oops:

please explain where I am going wrong.
Join the discussion

by Stuart@KaplanGMAT » Fri Sep 18, 2009 8:23 pm
uptowngirl92 wrote:P(Mike winning)=P(Ben looses). P(Mike wins) . P(Rob looses)
P(Rob winning)=P(Ben looses). P(Mike looses) . P(Rob wins)

Adding the two will give us the answer right?

Now given,P(mike wins)=1/4.Therefore,P(mike looses)=3/4
Similarly,P(rob wins)=1/3.Therefore,P(rob looses)=2/3
And,P(ben wins)=1/7.Therefore,P(ben looses)=6/7

Putting the values,
6/7.1/4.2/3 + 6/7.3/4.1/3 = 1/7+3/14=5/14 :oops:

please explain where I am going wrong.
You're double counting losses.

The only conditional part is Ben losing. There's no reason to multiply your individual events by the probability of Rob/Mike losing as well.

So, if you just want to add the two events:

Ben losing then Rob winning = 6/7 * 1/3 = 6/21 = 2/7

Ben losing then Mike winning = 6/7 * 1/4 = 6/28 = 3/14

2/7 + 3/14 = 4/14 + 3/14 = 7/14 = 1/2
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion